# Beyond Ensemble Averages: Leveraging Climate Model Ensembles for Subseasonal Forecasting

### v3

@@ 5.2 Baselines
Finally, we consider a baseline of a linear regression model applied to input features corresponding to ensemble member predictions: $\mathbf{x}_{t,l}=[u^{(1)}_{t,l},\dots,u^{(K)}_{t,l}]$. Then the model’s output
@@ 5.3 Learning-based methods
In contrast to the linear regression baseline, here other climate variables are added to the input features: $\mathbf{x}_{t,l}=[u^{(1)}_{t,l},...,u^{(K)}_{t,l},v^{(1)}_{t,l},...,v^{(P)}_{t,l}]$. Then the model’s output is defined with Eq. 8. Because the feature vector is higher dimensional here than for the baseline, the learned $\theta_{l}$ is also higher dimensional. We train a different model for each spatial location. In our experiments with linear models, we do not include positional encoding $(\mathbf{z}_{l}^{(1)},\mathbf{z}_{l}^{(2)})$ as input features, since they would be constants for that location’s linear model.
@@
In the context of quantile regression, we train a random forest quantile regressor (RFQR, Meinshausen [45]), which grows trees the same way as the original random forest while storing all training samples. To make a prediction for a test point, the RFQR computes a weight for each training sample that corresponds to the number of leafs (across all trees in the forest) that contain the test sample and the training sample. The RFQR prediction is then a quantile of the weighted training samples across all leafs that contain the test sample. We show a figure representation of the RFQR in Appendix D.2. With this formulation, training a single RFQR for all locations is computationally demanding, so we train individual RFQRs for every location.
The RFs are often referred to as the best off-the-shelf classifiers [46] even using the default hyperparameters [47]. Our cross-validation (CV) and grid search experiments show that the RFs hyperparameters have little impact on the accuracy, and thus we use the default parameters for RFs from the Scikit-learn library [48].
@@
Note that here we use capital letters because the input to our U-Net consists 2-D spatial maps, which are represented as matrices instead of vectors. The model output is a spatial map of the predicted target. This process is illustrated in Figure 1(c).
For the U-Net, we modify an available PyTorch implementation [50]. The training set consists of $249$ samples (images), which may be considered relatively limited for CNN training. To address this concern, we conduct bootstrapping experiments for the U-Net architecture, offering detailed insights into the impact of sample size on model performance. Further details are presented in Appendix C.3. We use 10-fold CV over our training data and grid search to select parameters such as learning rate, weight decay, batch size, and number of epochs. The Adam optimizer [51] is used in all experiments. After selecting hyperparameters, we train the U-Net model with those parameters on the full training dataset. The validation set is used to perform feature importance analysis. For regression, we train using squared error loss. In the context of quantile regression, we initialize the weights with those learned on squared error loss and then train on the quantile loss Eq. 9.
@@
Model stacking can improve model performance by combining the outputs of several models (usually called base models) [52]. In our case, linear regression, random forests, and the U-Net are substantially different in terms of architecture and computation, and we observe that they produce qualitatively different forecasts. We stack the linear model, random forest, and U-Net forecasts using a nonparametric approach:
@@
where $h$ is a simple feed-forward neural network with a non-linear activation, $\hat{y}_{t,l}^{\text{LR}},\hat{y}_{t,l}^{\text{RF}},\hat{y}_{t,l}^{\text{UNET}}$ are the predictions of a linear model, random forest and the U-Net correspondingly and referred to as ”base models”. One stacking model is trained to make predictions for any spatial location. Figure 1(b) demonstrates the stacking model’s input format and framework: the features are predictions from other ML models, and there is no spatial information. Model stacking can improve the forecast quality by combining predictions from three forecasting paradigms – spatial independence, conditional spatial independence, and spatial dependence (Section 5.4), and is analogous to the multi-model ensemble approach commonly used in weather and climate forecasting. The architecture details can be found in Appendix D.
For model stacking, we apply the following procedure: the base models are first trained on half of the training data, and predicted values on the second half are used to train the stacking model $h$. Then we retrain the base models on all the training data and apply the trained stacked model to the outputs of the base models. The proposed procedure helps avoid overfitting.
@@ 5.4 Models of spatial variation
Figure 1: An illustration of different forecasting paradigms: (a) spatial independence models with a model for each spatial location, no accounting for spatial information; (b) conditional spatial independence models with one model for all locations, might consider the spatial information; (c) spatial dependence models that account for the spatial information by the design. For temperature prediction, we replace “precipitation” in the illustration with “temperature”, but the overall structure remains the same.
@@ 6 Experimental setup
In this section, we provide details on the experimental setup including positional encoding, removing climatology, and evaluation metrics. Data preprocessing details are presented in Appendix E.
@@ 6.1 Positional encoding and inputs
Positional encoding [53] is a technique used in natural language processing (NLP) to inject positional information into data. In sequence-based tasks, such as language translation or text generation, the order of elements in the input sequence is important, but neural networks do not naturally capture this information. PE assigns unique encodings to each position in the sequence, which are then added to the original input before being processed by the model. This enables the model to consider the order and relative positions of elements, improving its ability to capture local and global context within a sequence and make accurate predictions [54, 55, 56]. This technique is helpful to represent the positional information outside the NLP tasks [57, 58, 59]. Several of our models use the spatial location as an input feature. Rather than directly using latitudes and longitudes, we use PE [53]:
@@
where $l$ is a longitude or latitude value, $d=12$ is the dimensionality of the positional encoding, and $i\in\{1,\dots,d\}$ is the index of the positional encoding vector. For the U-Net model, PE vectors are transformed into images in the following way: we take every value in the vector and fill the image of desired size with this value. So, there are $d$ images with the corresponding PE values. For the RF models, PE vectors can be used as they are.
@@
    four climate variables: relative humidity, pressure, geopotential height, and temperature (if target is precipitation) or precipitation (if target is temperature) two months before the target month,
@@
SSTs are usually represented as eight PCs, the embedding vector size is usually $d=12$ as we describe Section 6.1. For example, using the NCEP-CFSv2 members, there are $\underbrace{24}_{K}+\underbrace{4+5+8}_{P}+\underbrace{12}_{d}\times 2=65$ input features for every time step and location.
@@ 6.2 Evaluation metrics
Then the average $R^{2}$ for all locations is calculated as
@@
We also report the standard error (SE), median, and 90th percentile of $\{{\rm MSE}_{l}\}_{l}$. We say that the difference between two models is significant if their MSE $\pm$ SE intervals do not overlap. Note that the standard errors provided here should be used with caution since there are significant spatial correlations in the MSE values across locations, so we do not truly have $L$ independent samples from an asymptotically normal distribution.
@@
For the quantile regression task, we report mean quantile loss (9) across all locations.
@@ 7 Experimental results
In this section, we report the predictive skill of different models applied to SSF over the continental U.S.  using NCEP-CFSv2 ensemble members for regression and quantile regression. Precipitation forecasting is known to be more challenging compared to temperature forecasting [60]. The results for the NASA-GMAO dataset are presented in Appendix A. The skill of different models on the tercile classification task is presented in Appendix B for both datasets. Recall that all methods are trained on data spanning January 1985–September 2005, with data spanning October 2005–December 2010 used for validation (i.e., model selection and hyperparameter tuning). Test data spanning 2011 to 2020 was not viewed at any point of the model development and training process, and only used to evaluate the predictive skill of our trained models on previously unseen data; we refer to this period as the “test period”. The code is available at https://github.com/elena-orlova/SSF-project.As a navigation tool for the reader, Table 2 gives references to the presented results for different tasks.
@@ 7.1 Regression
Precipitation regression results are presented in Table 3. While the individual ML approaches produce results generally similar to those of the baselines, the stacked ML model, in particular, outperforms the baseline models in almost all metrics. Note that the best $R^{2}$ value, associated with the stacked model, is still near zero; while this is a significant improvement over, for example, the ensemble mean, which has an $R^{2}$ value of -0.08, the low values for all methods indicate the difficulty of the forecasting problem. It is important to note that $R^{2}$ measures the accuracy of a model relative to a baseline corresponding to the mean of the target over the test period – that is, relative to a model that could never be used in practice as a forecaster because it uses future observations. The best practical analog to this would be the mean of the target over the train period – what we call the “historical mean” or climatology model. These two models are not the same, possibly because of the nonstationarity of the climate [34]. Thus, even when our $R^{2}$ values are negative (i.e., we perform worse than the impractical mean of the target over the test period), we still perform much better than the practical climatology predictor. The model stacking approach is applied to the models that are trained on all available features (i.e., ensemble members, PE, climate variables; linear regression is trained on all features except PE). We decide what models to include in the stacking approach based on their performance on validation data. The low 90th percentile error implies that our methods not only have high skill on average, but also that there are relatively few locations with large errors. While acknowledging the overall performance may not be exceptional, it is important to recognize the potential of machine learning methods in improving the quality of estimates relative to the standard baselines. To further evaluate the capabilities of the stacking approach, we apply the stacking model to the baseline predictions, which include historical and ensemble means, as well as linear regression. The performance of the stacked baseline model exceeds that of any of the individual baseline models and is similar to the performance of the stacked ML approach in terms of the R2 metric. However, the stacked ML approach outperforms it in all MSE-based metrics, indicating that the ML techniques can still provide additional skill even for as notoriously challenging a quantity as precipitation.
Table 3: Results for precipitation regression the using NCEP-CFSv2 ensemble, with errors reported over the test period. LR refers to linear regression on all features including ensemble members, lagged data, climate variables and SSTs. Model stacking is performed on models that are trained on all features. The best results are in bold. MSE is reported in mm.
@@
Figure 2 illustrates performance of key methods with $R^{2}$ heatmaps over the U.S. to highlight spatial variation in errors. The RF and U-Net $R^{2}$ fields are qualitatively similar, but they are still quite different in certain states such as Georgia, North Carolina, Virginia, Utah, and Colorado. The LR map is noticeably poor across most of the regions. The stacked model’s heatmap reveals large regions where its predictive skill exceeds that of all other methods. Note that model stacking yields relatively accurate predictions even in regions where the three constituent models individually perform poorly (e.g., southwestern Arizona), highlighting the generalization abilities of our stacking approach. All methods tend to have higher accuracy on the Pacific Coast, in Midwest, and in southern states such as Alabama and Missouri. The heatmaps of the stacking model outputs look similar. The stacking model applied to the baselines has better $R^{2}$ scores in California compared to the stacked ML methods. However, the stacking model on ML models reveals larger positive $R^{2}$ regions and fewer dark red spots, particularly evident in New Mexico, Minnesota, and Utah.
@@
Table 4 shows results for 2-meter temperature regression. The learning-based models, especially the random forest and stacked model, significantly outperform the baseline models in terms of MSE and $R^{2}$ score. The random forest also outperforms linear regression and the U-Net. Note that LR, U-Net, and RF are trained without using SSTs information, since SSTs features yielded worse performance over the validation period. Figure 3 illustrates the performance of these methods with $R^{2}$ heatmaps over the U.S. As expected, the model stacking approach shows the best results across spatial locations. We notice that there are still regions such as the West, some regions in Texas, Florida and Georgia where all models tend to achieve a negative $R^{2}$ score.
Table 4: Results for temperature regression using the NCEP-CFSv2 ensemble, with errors reported over the test period. LR refers to linear regression on all features including ensemble members, lagged data, land variables. Model stacking is performed on models that are trained on all features except SSTs. The best results are in bold. MSE is reported in ${}^{\circ}C$.
@@ 7.2 Quantile regression
We explore the use of quantile regression to predict values $z$ so that “there’s a 90% chance that the average temperature will be below $z^{\circ}$ at your location next month” – or, equivalently, “there is a 10% chance that the average temperature will exceed $z^{\circ}$ at your location next month.” In this sense, quantile regression focused on the $90$-th percentile predicts temperature and precipitation extremes, a task highly relevant to a myriad of stakeholders. We train a linear regression model fitting the quantile loss (Linear QR), a random forest quantile regressor (RFQR), [45], a U-Net, and the stacked model. Details of the Linear QR and the RFQR are discussed in Section 5. The below experimental results show that temperature extremes can be predicted with high accuracy by the learning-based models (particularly our stacked model), in stark contrast to historical quantiles or ensemble quantiles in the case of temperature quantile regression. The results for precipitation are less striking overall, though the learned models are significantly more predictive in some locations as for the precipitation quantile regression task.
@@
For each location, the 90th percentile value is calculated based on the historical data. For the ensemble 90th percentile, we simply take the 90th percentile of the $K$ ensemble members. Table 5 summarizes results for precipitation quantile regression using the NCEP-CFSv2 ensemble. Our stacked model is able to significantly outperform all baselines. The performance illustration is given in the Appendix A, Figure 10.
@@
Table 6 summarizes results for temperature quantile regression using the NCEP-CFSv2 ensemble. Note that we do not include SST features for temperature quantile regression in our learned models. We observe that all of our learned models are able to significantly outperform all baselines. In Figure 4, we show the heatmaps of quantile loss of baselines and our learned models. We observe that the learned models produce different quality predictions, and the stacked model is able to pick up useful information from them. For example, in Arizona and Texas, the Linear QR, U-Net, and RFQR show some errors but in different locations, and the stacked model is able to exploit the advantages of each model.
Table 6: Test results for temperature quantile regression using NCEP-CFSv2 dataset, with target quantile = 0.9. Linear QR refers to a linear quantile regressor. RFQR corresponds to a Random Forest Quantile Regressor. Model stacking is performed on models that are trained on all features except for SSTs. Learned models are able to predict highly likely temperature ranges accurately, meaning there are fewer unpredicted temperature spikes. The best results are in bold.

### v4

@@ 5.2 Baselines
Finally, we consider, as a baseline, a linear regression model applied to input features corresponding to ensemble member predictions: $\mathbf{x}_{t,l}=[u^{(1)}_{t,l},\dots,u^{(K)}_{t,l}]$. Then, the model’s output
@@ 5.3 Learning-based methods
In contrast to the linear regression baseline, here other climate variables are added to the input features: $\mathbf{x}_{t,l}=[u^{(1)}_{t,l},...,u^{(K)}_{t,l},v^{(1)}_{t,l},...,v^{(P)}_{t,l}]$. Then the model’s output is defined with Eq. 10. Because the feature vector is higher dimensional here than for the baseline, the learned $\theta_{l}$ is also higher dimensional. We train a different model for each spatial location. In our experiments with linear models, we do not include positional encoding $(\mathbf{z}_{l}^{(1)},\mathbf{z}_{l}^{(2)})$ as input features, since they would be constants for each location’s linear model.
@@
In the context of quantile regression, we train a random forest quantile regressor (RFQR, Meinshausen [45]), which grows trees the same way as the original random forest while storing all training samples. To make a prediction for a test point, the RFQR computes a weight for each training sample that corresponds to the number of leaves (across all trees in the forest) that contain the test sample and the training sample. The RFQR prediction is then a quantile of the weighted training samples across all leaves that contain the test sample. We show a figure representation of the RFQR in Section D.2. With this formulation, training a single RFQR for all locations is computationally demanding, so we train individual RFQRs for every location.
Random forests are often referred as the best off-the-shelf classifiers [46] even using the default hyperparameters [47]. Our cross-validation (CV) and grid search experiments show that the RFs hyperparameters have little impact on the accuracy, and thus, we use the default parameters for RFs from the Scikit-learn library [48].
@@
Note that here, we use capital letters because the input to our U-Net consists of 2-D spatial maps, which are represented as matrices instead of vectors. The model output is a spatial map of the predicted target. This process is illustrated in Figure 1(c).
For the U-Net, we modify an available PyTorch implementation [50]. The training set consists of $249$ samples (images), which may be considered relatively limited for CNN training. To address this concern, we conduct bootstrapping experiments for the U-Net architecture, offering detailed insights into the impact of sample size on model performance. Further details are presented in Appendix C.2. We use a 10-fold CV over our training data and grid search to select parameters such as learning rate, weight decay, batch size, and number of epochs. The Adam optimizer [51] is used in all experiments. After selecting hyperparameters, we train the U-Net model with those parameters on the full training dataset. The validation set is used to perform feature importance analysis. For regression, we train using squared error loss. In the context of quantile regression, we initialize the weights with those learned on squared error loss and then train on the quantile loss Eq. 11.
@@
Model stacking can improve model performance by combining the outputs of several models (usually called base models) [52]. In our case, linear regression, random forests, and the U-Net are substantially different in architecture and computation, and we observe that they produce qualitatively different forecasts. We stack the linear model, random forest, and U-Net forecasts using a nonparametric approach:
@@
where $h$ is a simple feed-forward neural network with a non-linear activation and $\hat{y}_{t,l}^{\text{LR}},\hat{y}_{t,l}^{\text{RF}},\hat{y}_{t,l}^{\text{UNET}}$ are the predictions of a linear model, random forest and the U-Net correspondingly and referred to as “base models”. One stacking model is trained to make predictions for any spatial location. Figure 1(b) with input features that are predictions from other ML models and no PE vectors demonstrates the stacking model’s framework. Model stacking can improve the forecast quality by combining predictions from three forecasting paradigms – spatial independence, conditional spatial independence, and spatial dependence (Section 5.4), and is analogous to the multi-model ensemble approach commonly used in weather and climate forecasting. The architecture details can be found in Appendix D.
We apply the following procedure for model stacking: the base models are first trained on half of the training data, and predicted values on the second half are used to train the stacking model $h$. Then, we retrain the base models on all the training data and apply the trained stacked model to the outputs of the base models. The proposed procedure helps to avoid overfitting.
@@ 5.4 Models of spatial variation
Figure 1: An illustration of different forecasting paradigms: (a) spatial independence models with a model for each spatial location, no accounting for spatial information; (b) conditional spatial independence models with one model for all locations, might consider the spatial information; (c) spatial dependence models that account for the spatial information by design. We replace “precipitation” in the illustration with “temperature” for temperature prediction, but the overall structure remains the same.
@@ 6 Experimental setup
This section provides details on the experimental setup, including positional encoding, removing climatology, and evaluation metrics. Data preprocessing details are presented in Appendix E.
@@ 6.1 Positional encoding and inputs
Positional encoding [53] is a technique used in natural language processing (NLP) to inject positional information into data. In sequence-based tasks, such as language translation or text generation, the order of elements in the input sequence is important, but neural networks do not naturally capture this information. PE assigns unique encodings to each position in the sequence, which are then added to the original input before being processed by the model. This enables the model to consider the order and relative positions of elements, improving its ability to capture local and global context within a sequence and make accurate predictions [54, 55, 56]. This technique is helpful to represent the positional information outside the original NLP tasks [57, 58, 59]. Several of our models use the spatial location as an input feature. Rather than directly using latitudes and longitudes, we use PE [53]:
@@
where $l$ is a longitude or latitude value, $d=12$ is the dimensionality of the positional encoding, and $i\in\{1,\dots,d\}$ is the index of the positional encoding vector. For the U-Net model, PE vectors are transformed into images in the following way: we take every value in the vector and fill the image of the desired size with this value. So, there are $d$ images with the corresponding PE values. For the RF models, PE vectors can be used as they are.
6.2 Models’ inputs details
@@
    four climate variables: relative humidity, pressure, geopotential height, and temperature (if the target is precipitation) or precipitation (if the target is temperature) two months before the target month,
@@
SSTs are usually represented as eight PCs, and the embedding vector size is usually $d=12$ as we describe Section 6.1. For example, using the NCEP-CFSv2 members, there are $\underbrace{24}_{K}+\underbrace{4+5+8}_{P}+\underbrace{12}_{d}\times 2=65$ input features for every time step and location. Figure 1 provides an illustration of these input features.
@@ 6.3 Evaluation metrics
Then, the average $R^{2}$ for all locations is calculated as
@@
We also report the standard error (SE), median, and 90th percentile of $\{{\rm MSE}_{l}\}_{l}$. We say the difference between the two models is significant if their MSE $\pm$ SE intervals do not overlap. Note that the standard errors provided here should be used with caution since there are significant spatial correlations in the MSE values across locations, so we do not truly have $L$ independent samples from an asymptotically normal distribution.
@@
For the quantile regression task, we report mean quantile loss from Eq. 11 across all locations.
@@ 7 Experimental results
In this section, we report the predictive skill of different models applied to SSF over the continental U.S.  using NCEP-CFSv2 ensemble members for regression and quantile regression. Precipitation forecasting is known to be more challenging compared to temperature forecasting [60]. The results for the NASA-GMAO dataset are presented in Appendix A. The skill of different models on the tercile classification task is presented in Appendix B for both datasets. Recall that all methods are trained on data spanning January 1985–September 2005, with data spanning October 2005 - December 2010 used for validation (i.e., model selection and hyperparameter tuning). Test data spanning 2011 to 2020 was not viewed at any point of the model development and training process and only used to evaluate the predictive skill of our trained models on previously unseen data; we refer to this period as the “test period”. As a navigation tool for the reader, Table 2 gives references to the presented results for different tasks.
@@ 7.1 Regression
Precipitation regression results are presented in Table 3. While the individual ML approaches produce results generally similar to those of the baselines, the stacked ML model, in particular, outperforms the baseline models in almost all metrics. Note that the best $R^{2}$ value, associated with the stacked model, is still near zero; while this is a significant improvement over, for example, the ensemble mean, which has an $R^{2}$ value of -0.08, the low values for all methods indicate the difficulty of the forecasting problem. It is important to note that $R^{2}$ measures the accuracy of a model relative to a baseline corresponding to the mean of the target over the test period – that is, relative to a model that could never be used in practice as a forecaster because it uses future observations. The best practical analog to this would be the mean of the target over the train period – what we call the “historical mean” or climatology model. These two models are not the same, possibly because of the nonstationarity of the climate [34]. Thus, even when our $R^{2}$ values are negative (i.e., we perform worse than the impractical mean of the target over the test period), we still perform much better than the practical climatology predictor. The model stacking approach is applied to the models trained on all available features (i.e., ensemble members, PE, climate variables; linear regression is trained on all features except PE). We decide what models to include in the stacking approach based on their performance on validation data. The low 90th percentile error implies that our methods not only have high skill on average but also that there are relatively few locations with large errors. While acknowledging the overall performance may not be exceptional, it is important to recognize the potential of machine learning methods in improving the quality of estimates relative to the standard baselines. To further evaluate the capabilities of the stacking approach, we also apply the approach to the baseline predictions, which include historical and ensemble means, as well as linear regression. The performance of the stacked baseline model exceeds that of any of the individual baseline models and is similar to the performance of the stacked ML approach in terms of the $R^{2}$ metric. However, the stacked ML approach outperforms it in all MSE-based metrics, indicating that the ML techniques can still provide additional skill even for as notoriously challenging a quantity as precipitation.
Table 3: Results for precipitation regression the using NCEP-CFSv2 ensemble, with errors reported over the test period. LR refers to linear regression on all features, including ensemble members, lagged data, climate variables, and SSTs. ML model stacking is performed on models that are trained on all features. The best results are in bold. MSE is reported in squared mm.
@@
Figure 2 illustrates performance of key methods with $R^{2}$ heatmaps over the U.S. to highlight spatial variation in errors. The RF and U-Net $R^{2}$ fields are qualitatively similar, but they are still quite different in certain states such as Georgia, North Carolina, Virginia, Utah, and Colorado. The LR map is noticeably poor across most of the regions. The stacked ML model’s heatmap reveals large regions where its predictive skill exceeds that of all other methods. Note that model stacking yields relatively accurate predictions even in regions where the three constituent models individually perform poorly (e.g., southwestern Arizona), highlighting the generalization abilities of our stacking approach. All methods tend to have higher accuracy on the Pacific Coast, in the Midwest, and in southern states such as Alabama and Missouri. The stacking model heatmaps both look similar. The stacking model applied to the baselines has better $R^{2}$ scores in California compared to the stacked ML methods. However, the stacked ML model reveals larger positive $R^{2}$ regions and fewer dark red spots, particularly evident in New Mexico, Minnesota, and Utah.
@@
Table 4 shows results for 2-meter temperature regression. The learning-based models, especially the random forest and stacked model, significantly outperform the baseline models in terms of MSE and $R^{2}$ score. The random forest also outperforms linear regression and the U-Net. Note that LR, U-Net, and RF are trained without using SST information since SST features yielded worse performance over the validation period. Figure 3 illustrates the performance of these methods with $R^{2}$ heatmaps over the U.S. As expected, the model stacking approach shows the best results across spatial locations. We notice that there are still regions such as the West, some regions in Texas, Florida, and Georgia where all models tend to achieve a negative $R^{2}$ score.
Table 4: Results for temperature regression using the NCEP-CFSv2 ensemble, with errors reported over the test period. LR refers to linear regression on all features including ensemble members, lagged data, land variables. Model stacking is performed on models that are trained on all features except SSTs. The best results are in bold. MSE is reported in squared ${}^{\circ}C$.
@@ 7.2 Quantile regression
We explore the use of quantile regression to predict values $z$ so that “there’s a 90% chance that the average temperature will be below $z^{\circ}$ at your location next month” – or, equivalently, “there is a 10% chance that the average temperature will exceed $z^{\circ}$ at your location next month.” In this sense, quantile regression focused on the $90$-th percentile predicts temperature and precipitation extremes, a task highly relevant to many stakeholders. We train a linear regression model fitting the quantile loss (Linear QR), a random forest quantile regressor (RFQR), [45], a U-Net, and the stacked model. The Linear QR and the RFQR details are discussed in Section 5. The below experimental results show that temperature extremes can be predicted with high accuracy by the learning-based models (particularly our stacked model), in stark contrast to historical quantiles or ensemble quantiles in the case of temperature quantile regression. The results for precipitation are less striking overall, though the learned models are significantly more predictive in some locations on this quantile regression task.
@@
For each location, the 90th percentile value is calculated based on the historical data. For the ensemble 90th percentile, we simply take the 90th percentile of the $K$ ensemble members. Table 5 summarizes results for precipitation quantile regression using the NCEP-CFSv2 ensemble. Our stacked model is able to significantly outperform all baselines. The performance illustration is given in Appendix A, Figure 10.
@@
Table 6 summarizes results for temperature quantile regression using the NCEP-CFSv2 ensemble. Note that we do not include SST features for temperature quantile regression in our learned models. We observe that all of our learned models are able to significantly outperform all baselines. In Figure 4, we show the heatmaps of quantile loss of baselines and our learned models. We observe that the learned models produce predictions with varied quality, and the stacked model can pick up useful information from them. For example, in Arizona and Texas, the Linear QR, U-Net, and RFQR show some errors but in different locations, and the stacked model can exploit the advantages of each model.
Table 6: Test results for temperature quantile regression using NCEP-CFSv2 dataset, with target quantile = 0.9. Linear QR refers to a linear quantile regressor. RFQR corresponds to a Random Forest Quantile Regressor. Model stacking is performed on models that are trained on all features except for SSTs . Learned models can predict highly likely temperature ranges accurately, meaning there are fewer unpredicted temperature spikes. The best results are in bold.